# Leveraging Local Patch Alignment to Seam-cutting for Large Parallax Image Stitching

### v2

@@ IV-A Quantitative evaluation
Table I presents the quantitative evaluation results of different seam cutting methods before and after adding LPAM. Evidently, LPAM can substantially improve the seam quality of any seam cutting method across all large parallax image datasets. We also show some of the seam comparison results in Fig. 5, where the seams are shown using colormaps based on the quality error defined by Eq. (2). Through local alignment of misaligned patches, our method effectively improves the quality of seam cutting and the final visual stitching results.
Figure 5: Visual seam quality comparisons w/ and w/o LPAM. Top: initial seams w/o LPAM. Bottom: final seams w/ LPAM.
Our method improves the seam cutting by locally realigning the misaligned patches along the stitching seam. To verify its effectiveness, we also compare the global alignment accuracy of the overlapping region in different seam cutting methods, both before and after incorporating LPAM. The comparison results are shown in Table II. Low PSNR and SSIM values indicate that accurately aligning the input images is challenging. Our LPAM can continuously improve the alignment accuracy, regardless of the estimated stitching seam, thus generating visually superior stitching results.
TABLE I: Quantitative evaluation of seam cutting w/ and w/o LPAM on different image datasets.
@@
Baseline+LPAM 0.078 31.90 0.784 0.099 0.096 28.31 0.696 0.138 0.075 28.78 0.780 0.103
Perception+LPAM 0.074 31.23 0.795 0.084 0.103 25.37 0.682 0.127 0.072 28.82 0.801 0.084
Iterative+LPAM 0.076 30.87 0.800 0.087 0.103 25.37 0.682 0.127 0.072 28.91 0.786 0.101
Quaternion+LPAM 0.080 29.41 0.783 0.085 0.105 26.09 0.677 0.136 0.067 29.25 0.812 0.078
TABLE II: Quantitative evaluation of global alignment accuracy w/ and w/o LPAM on different image datasets.
Dataset Parallax [24] SEAGULL [25] Our dataset
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
REW [6] 13.539 0.630 13.669 0.551 14.928 0.682
Baseline+LPAM 13.591 0.636 13.739 0.560 14.966 0.685
Perception+LPAM 13.589 0.636 13.758 0.564 14.973 0.687
Iterative+LPAM 13.593 0.636 13.759 0.561 14.974 0.685
Quaternion+LPAM 13.602 0.636 13.739 0.561 15.009 0.689
IV-B Visual Comparison
@@
We comprehensively compare the visual stitching results of our method with those of methods integrating the seam cutting technique. These include Microsoft ICE [^2] https://www.microsoft.com/en-us/research/project/image-composite-editor, [18, 19, 20], Parallax [24], and SEAGULL [25]. Fig. 6 shows a particularly challenging example sourced from [24]. The input images and the result of Parallax are retrieved from the website [^3] http://web.cecs.pdx.edu/~fliu/project/stitch/. Notably, Microsoft ICE, the baseline method, and Iterative seam cutting [19] fail to produce satisfactory results (obvious artifacts are marked by green arrows). The simple integration of our LPAM improves the stitching quality and achieves visual results comparable to those of Parallax and SEAGULL. It should be noted that Parallax and SEAGULL need to estimate multiple seam cutting results from multiple alignment candidates and then select the optimal one. In contrast, our LPAM provides a more straightforward and effective solution for improving stitching quality.
(a) Input images
(b) Microsoft ICE
(c) Parallax [24]
(d) SEAGULL [25]
(e) Baseline
(f) Baseline+LPAM
(g) Perception [18]
(h) Perception+LPAM
(i) Iterative [19]
(j) Iterative+LPAM
(k) Quaternion [20]
(l) Quaternion+LPAM
@@
Figure 6: Comparisons of image stitching results via seam cutting methods.
IV-C Computational Efficiency
@@
We evaluate the computational efficiency of LPAM when integrated with other seam cutting methods. Experiments are conducted on three image datasets using a 3.6GHz Intel Core i7 CPU with 16GB memory. We record the elapsed time for each seam cutting method [^4] Time recorded is solely for seam cutting, excluding image alignment. and for LPAM individually. The results are presented in Table III. In LPAM, constructing dense correspondence dominates the elapsed time. This process is influenced by the number and resolution of extracted patches, which vary significantly across different image pairs and datasets. Notably, when applied to various seam cutting methods, LPAM’s processing time remains relatively stable, ranging from 6-10 seconds [^5] LPAM is currently implemented in MATLAB, integrating with C++ could further improve its efficiency..
@@
TABLE III: Elapsed time of applying LPAM to different seam cutting methods on the three image datasets.
@@
(a) Input images
(b) Initial stitching seam and result
(c) Stitching seam and result w/o LPAM
(d) Stitching seam and result w/ LPAM
IV-D Discussion and Limitation
Our method assumes that the stitching seam experiences partial failure such that only several patches along the seam need realignment. In our implementation, a seam is considered to be plausible if $\max(Q)\leq 1.5\cdot\mathrm{mean}(Q)$. Nevertheless, as shown in Fig. 7(a,b), this inequality can still hold even if most pixels along the seam fail. In such cases, our method fails since it regards the seam as plausible, and no patches are extracted for realignment. Designing a superior algorithm to identify the “problematic” pixels along the stitching seam could potentially resolve this issue. We leave this as future work.
Furthermore, in Eqs. (4,5), we introduce a modification to the vector flow between separated patches to smoothly merge them into pre-aligned results. However, when significant structural misalignment exists within small patches, this modification is insufficient to ensure smoothness, as shown in Fig. 7(c,d). The extracted patch (marked by the black box) is too small for effective smoothness modification. Consequently, our LPAM struggles to strike a balance between accurate alignment and smooth merging. Enlarging the patch size could yield smoother merging. We leave this as future work.
V Conclusion
In this paper, we propose a simple local patch alignment module that can effectively and efficiently improve the stitching performance of the seam cutting methods for large parallax image stitching. Firstly, we evaluate the pixels along the initial seam and extract misaligned components of the seam. For each such component, we separate the enclosing rectangular patches in the pre-aligned images and construct modified dense correspondence using SIFT flow to align these patches. Finally, we calculate a local seam and merge it into the initial seam. Experimental results on various stitching datasets demonstrate the superiority of our method.
Acknowledgment
This work is partially supported by the Natural Science Foundation of Henan Province under Grant 222300420140.

### v3

@@ 4.1 Quantitative evaluation
4.1.1 Evaluation across different seam-cutting methods
Table 1 presents the quantitative evaluation results of adding LPAM to different seam-cutting methods under the same alignment results via REW [16]. Evidently, LPAM can substantially improve the seam quality across different seam-cutting methods on all large parallax stitching datasets. We also show part of the seam comparison results in Fig. 5, where the seams are shown using colormaps based on the quality error defined by Eq. (2). Through the local alignment of misaligned patches, our method effectively improves the quality of seam-cutting and the final visual stitching results.
4.1.2 Evaluation across different alignment methods
As shown in Table 2, we also evaluate the LPAM when applying it to a single seam-cutting method under the different representative alignment methods, including global homography (Homo), ANAP [21], GSP [6], REW [16] and the state-of-the-art learning-based method UDIS++ [25]. LPAM still achieves stable improvement of the seam quality across different alignment results.
Table 3: Ablation study of applying LPAM to basic Euclidian-based seam-cutting. Alignment results are obtained via REW [16].
@@ 4.1.2 Evaluation across different alignment methods
$\beta=\emptyset$ 0.069 31.38 0.819 0.082 0.084 28.24 0.768 0.101 0.069 29.33 0.815 0.086
$\beta=1$ 0.080 30.51 0.770 0.103 0.099 28.75 0.677 0.143 0.076 28.73 0.774 0.103
$\beta=2$ 0.080 30.49 0.771 0.103 0.098 28.88 0.680 0.142 0.076 28.73 0.774 0.105
$\beta=4$ 0.080 30.73 0.775 0.102 0.098 28.23 0.683 0.142 0.076 28.72 0.771 0.106
$\beta=8$ 0.078 31.85 0.782 0.100 0.096 28.84 0.697 0.137 0.075 28.76 0.778 0.104
$\beta=16$ 0.078 30.92 0.784 0.099 0.095 29.30 0.706 0.132 0.075 28.84 0.782 0.103
$7\times 7$ 0.084 29.90 0.762 0.110 0.099 28.05 0.679 0.145 0.076 28.66 0.774 0.107
$11\times 1$ 11 0.081 30.13 0.771 0.104 0.098 28.16 0.683 0.142 0.075 28.74 0.776 0.106
$31\times 31$ 0.078 31.69 0.783 0.099 0.098 27.52 0.692 0.137 0.076 28.78 0.784 0.102
$41\times 41$ 0.078 31.67 0.777 0.099 0.100 27.39 0.684 0.140 0.074 28.93 0.784 0.103
$51\times 51$ 0.081 31.80 0.775 0.100 0.099 28.05 0.684 0.144 0.075 28.91 0.782 0.103
4.2 Visual results comparison
@@
We comprehensively compare the visual stitching results of our method with those of methods integrating the seam-cutting technique. These include Microsoft ICE [^1] https://www.microsoft.com/en-us/research/project/image-composite-editor, [19, 20, 15], Parallax [31], and SEAGULL [22]. Fig. 7 shows a particularly challenging example sourced from [31]. The input images and the result of Parallax are retrieved from the website [^2] http://web.cecs.pdx.edu/~fliu/project/stitch/. Notably, Microsoft ICE, the baseline method, and Iterative seam-cutting [20] fail to produce satisfactory results (obvious artifacts are marked by green arrows). The simple integration of our LPAM improves the stitching quality and achieves visual results comparable to those of Parallax and SEAGULL. It should be noted that Parallax and SEAGULL need to estimate multiple seam-cutting results from multiple alignment candidates and select the optimal one. In contrast, our LPAM provides a more straightforward and effective solution for improving stitching quality. More visual comparisons are provided in the supplementary.
@@
Figure 7: Comparisons of image stitching results via different seam-cutting methods. (Best view in color and zoom in)
4.3 Ablation study
We conduct an ablation study to assess the performance of our LPAM under different parameter settings, as shown in Table 3. The first row denotes that we ablate the sigmoid function in Eq. (5) and use the original sift flow to align the patches. The 2nd to 6th rows represent that the $\beta$ in Eq. (4) is set to 1,2,4,8,16. The last five rows represent different patch sizes for seam quality evaluation in Eq. (2) (the patch size in Eqs. (6-9) is fixed at 21$\times$21 for fair comparison), including the “gray” row is performed using $21\times 21$ patch size. Comparison results confirm that our LPAM can achieve stable improvement under diverse parameter settings of $\beta$. The first row shows that ablating the sigmoid function improves the seam quality according to the seam quality metrics. However, it may introduce artifacts outside the separated rectangular patches, as shown in Fig. 3. The study also shows that our LPAM is relatively stable under different patch sizes. Small patch size may limit smooth merging, while a large one challenges alignment and costs more (we discuss this in the supplementary). Considering these, we set the patch size to $21\times 21$ to balance the smooth merging and computational efficiency.
4.4 Computational efficiency
@@
We evaluate the computational efficiency of LPAM when integrated with other seam-cutting methods. Experiments are conducted using a 3.6GHz Intel Core i7 CPU with 16GB of memory. We record the elapsed time for each seam-cutting method [^3] Time recorded is solely for seam-cutting, excluding initial image alignment. and for LPAM individually. The results are presented in Table 4. In LPAM, constructing dense correspondence dominates the elapsed time. This process is influenced by the number and resolution of extracted patches, which vary significantly across image pairs and datasets. Notably, when applied to various seam-cutting methods, LPAM’s processing time remains relatively stable, ranging from 6 to 10 seconds [^4] Currently, LPAM is implemented without integration with C++ for further acceleration.. The computational cost of each process in our LPAM is discussed in the supplementary.
@@
Table 4: Elapsed time of applying LPAM to different seam-cutting methods on the three image datasets.
@@
4.5 Discussion and limitation
In Eqs. (4,5), we modify the vector flow estimated by SIFT flow to ensure the aligned patch is smoothly consistent with pre-aligned images. However, when significant structural misalignment exists within small patches, this modification is insufficient to ensure smoothness, as shown in Fig. 6(c,d). The extracted patch (marked by the black box) is too small for effective smoothness modification. Consequently, our LPAM struggles to attain a harmonious balance between accurate patch alignment and smooth consistency. Adaptively enlarging the patch size could yield smoother consistency. We leave this as future work.
5 Conclusion
In this paper, we propose a simple LPAM that can effectively and efficiently enhance the stitching performance of seam-cutting methods for large parallax image stitching. Our method is founded on a crucial discovery: the seam-cutting method is a downstream process that depends on successful image alignment. We introduce an alignment-compensation paradigm to dissociate seam quality from initial image alignment. Experimental results of applying our LPAM to the state-of-the-art seam-cutting methods with diverse alignment results on various stitching datasets demonstrate the superiority and robustness of our proposed method.
Acknowledgments
This work is partially supported by the Natural Science Foundation of Henan Province, China under Grant 222300420140.